OpenStudy (anonymous):

MEDAL AND FAN HELP! The equation for line m is y=1/2x-5 Use this information and the following description to write an equation for line n, line p, and line r. Hint: y = mx + b is the slope intercept form. (a) Line m and line n have no solutions. Write an equation that can represent line n. (Hint: the equation must have a slope of ½ and any y-intercept except -5) (b) Line m and line p have infinitely many solutions. Write an equation that can represent line p. ( : your equation must be equivalent to y=1/2x – 5 so it can be a different form (with five on the left side))

OpenStudy (anonymous):

or you can multiply both sides of the equation by a number) (c) Line m and line r have one solution. Write an equation that can represent line r. (Hint: it can have any y-intercept and any slope except ½)

OpenStudy (princeharryyy):

a) y = 1/2x + t where t belongs to any real number

OpenStudy (anonymous):

for c it can have any y intercept and slope except 1/2

OpenStudy (princeharryyy):

for b can u tell me wat r they really trying to ask in hint

OpenStudy (princeharryyy):

@adelie.jade can u so, that i can help u

OpenStudy (anonymous):

(Hint: your equation must be equivalent to y=1/2x – 5 so it can be a different form (with five on the left side), or you can multiply both sides of the equation by a number)

OpenStudy (anonymous):

y=1/2x -5

OpenStudy (anonymous):

thats what the question marks are

OpenStudy (princeharryyy):

equations with same slopes never meet ... that means lines with same slope have no solutions they are asking for infinite solutions do u know this

OpenStudy (princeharryyy):

no matter what the intercept is whether 5 on left or right but the slopes got to be different for the two lines for the infinite solutions

OpenStudy (princeharryyy):

y=x-5 is one valid equation

OpenStudy (princeharryyy):

there can b others as well...

OpenStudy (anonymous):

ok thxs